OpenStudy (anonymous):

Algebra 2 Evaluate the expression below when x=3. 5x^2-12/x+3 A. 3 B. 11/3 C. 11/2 D. 26/3

OpenStudy (anonymous):

5*9 - 4 + 3=44 So, the answer must be none of the above

OpenStudy (anonymous):

How? I only have 4 choose

Vocaloid (vocaloid):

perhaps you're missing some parentheses in your equation? (5x^2-12)/(x+3) = (5(9)-12)/(6) = 33/6 = 11/3

OpenStudy (anonymous):

Thank you so much :) @Vocaloid